# A Survey on Quantum Machine Learning: Basics, Current Trends, Challenges, Opportunities, and the Road Ahead

### v2

@@ V-E Quantum Datasets Collection
Table II shows the collection of quantum datasets with key features. The following list provides a brief description of each dataset.
@@
  • The QDataSet [118], released in 2022, is composed of $52$ datasets based on simulations of one-qubit and two-qubit systems evolving in the presence or absence of noise. It provides a large-scale set of datasets for QML practitioners to train, benchmark, and develop classical and quantum algorithms for common tasks in quantum sciences, such as quantum control, quantum tomography, and noise spectroscopy. It has been generated using customized code running on base-level Python packages to facilitate interoperability and portability across common QML platforms. Each dataset consists of $10,000$ samples and includes a range of information (stored in list, matrix, or tensor format) regarding quantum systems and their evolution, such as quantum state vectors, drift and control Hamiltonians, unitaries, Pauli measurement distributions, time series data, pulse sequence data for square and Gaussian pulses, noise and distortion data. The total compressed size of the QDataSet (compressed with Pickle and zip) is around $14$ TB, while its uncompressed size is around $100$ TB.
@@
  • The QM7, QM8, and QM9 datasets are subsets of the GDB-13 [127] and GDB-17 [128] databases, which contain billions of organic molecules. The QM7 [119] is composed of $7,165$ molecules of up to $23$ atoms (including $7$ heavy atoms such as C, N, O, and S), represented using the Coulomb matrix. An extension of the QM7 dataset for multitask learning, called QM7B [120], has $13$ additional properties (e.g., polarizability, HOMO and LUMO eigenvalues, excitation energies) and contains a total of $7,211$ molecules. The QMB8 [121] has a training set of $10,000$ molecules with up to $8$ heavy atoms. The QM9 [122] contains $133,885$ molecules with up to $9$ heavy atoms. These databases serve as benchmarks for hybrid QML models and systematic identification of structure-property relationships.
@@
  • The PennyLane Datasets [48] are composed of the Quantum Chemistry (QChem) Datasets for common molecular systems and the Quantum Many-Body Physics (QMBP) Datasets. The QChem datasets contain the electronic structure data for $42$ different geometries of molecules such as linear hydrogen chains, metallic and non-metallic hydrides, and charged species. For each geometry of the molecule, it is possible to extract molecular data (i.e., information regarding the molecule), Hamiltonian data (i.e., Hamiltonian for the molecular system), tapering data (i.e., features based on Z2 symmetries of the molecular Hamiltonian for performing tapering), tapered observables data, and variational data. The QMBP datasets contain spin models displaying quantum correlations. For each spin system, datasets are generated for $1$-D lattices (linear chain) and $2$-D lattices (rectangular grid). Each dataset contains values for $100$ different configurations of tunable parameters such as the external magnetic field and coupling constants.
@@
  • The TensorFlow Quantum Data builds upon classical datasets such as MNIST [129] and Fashion-MNIST [130]. It implements the method proposed by Farhi et al. [123] to use binary encoding for representing each pixel with a qubit, with the state depending on the pixel’s value. According to the experimental results of Huang et al. [131], this method achieves quantum speedup in the fault-tolerant regime compared to the classical system.
@@
  • The Mendley Datasets for Quantum Circuits Mapping [124] helps to solve the circuit mapping problem as a classification task. Each dataset contains features related to the calibration data of the physical device and others related to the generated quantum circuit for specific IBM quantum machines, such as IBMQ Santiago, IBMQ Athens, and IBMQ 16 Melbourne.
@@
  • The QFlow Lite dataset [125] contains a Python-based software suite to train neural networks to recognize the state of a device and differentiate between states in experimental data. It consists of $1,001$ idealized simulated measurements with gate configurations sampling (stored as $100\times 100$ pixel maps) over different realizations of the same type of device. The QFlow Lite represents a reference dataset for researchers to use in their experiments for developing QML approaches and concepts. The expanded dataset, denominated QFlow 2.0, consists of $1,599$ idealized simulated measurements stored as $250\times 250$ pixel maps. In addition, the QFlow 2.0 dataset includes two sets of noisy simulated measurements, one with a noise level varied around $1.5$ times the optimized noise level, and the other with a noise level ranging from 00 to $7$ times the optimized noise level.
@@
  • The NTangled Datasets [126] consist of trained weights for three hardware-efficient Ansatzes), strongly entangling, and convolutional, varying the number of qubits, depths, and types of multipartite entanglement. The states are generated by sending product state inputs into PQCs. The NTangled Datasets are used to benchmark QML models for supervised learning classification tasks.
VI Tools and Technologies
@@
VI-A Quantum Computing Models
@@
The first concept of a quantum computer was introduced in 1982 by Richard Feynman [31]. Since then, a significant amount of research efforts have been conducted that resulted in the development of real quantum computers at our disposal for experiments [132]. Due to their different use cases and architectures, quantum computers can be classified into quantum gate models (or quantum circuits) and quantum annealers.
VI-A1 Quantum Gate Model
@@
The gate model involves operations in quantum computation using matrix operators, which are applied to the input qubits. These operators process the information and generate the final output distributions. They are designed with unitary constraints to qualify as valid quantum operators, making them reversible. The reversibility property is necessary for introducing some form of determinism in the quantum computation. A single operator is referred to as a gate in the quantum circuit and can be applied independently or grouped together with other similar quantum operators to perform computation on qubits. A gate can be combined with other gates in varying orientations and series such that they solve certain problems. The collection of gates arranged in configurations respective to the problem being solved along with qubits and measurements is referred to as a circuit (see Section II-D). Complex problems in the quantum domain can be solved by configuring gates in meaningful ways. The computation done with this quantum model is referred to as the gate model. Since it is easily interpretative and intuitive, it is nowadays the most used quantum computation model.
VI-A2 Quantum Annealer
@@
Quantum Annealing (QA) is an adiabatic quantum computation method that finds optimal solutions to problems. It uses quantum tunneling, entanglement, and superposition to find the solution in terms of the system’s energy. It usually corresponds to finding the global energy minimum in the quantum energy state plane. Instead of using temperature to explore the problem space, QA uses the laws of quantum mechanics to measure the energy state. A quantum annealer is relatively more noise-tolerant compared to the gate model. Quantum annealers shine when they are used to search for the best from many possible combinations. Such challenges are called combinatorial optimization problems (see Section IV-C4).
VI-B Quantum Hardware
@@
Single qubits can be implemented using different materials and principles. As shown in Figure 31, companies do not use the same technology for implementing their quantum computers. For example, IBM Quantum uses niobium and aluminum on a silicon base to achieve superconductivity under near absolute zero temperature [69]. IonQ uses trapped ion quantum made of Ytterbium, a rare metal [133]. Another example is Xanadu, which applies photonics (light particles) in the quantum processing unit [134]. Using ion and light has advantages over the superconductivity materials as it works under room temperature. The superconducting qubits are more scalable and are implemented by other companies like Google [63] and D-Wave [135]. It is exciting to see future innovations in this area. In this section, after discussing the set of conditions necessary for designing qubits (denoted as DiVincenzo’s criteria), we review the qubit technologies to understand their advantages and disadvantages and the reasons for their adoption in the industry. A detailed view of the quantum hardware technologies is also illustrated in Table III.
@@
TABLE III: Quantum hardware technologies with key features.
@@
IBM Superconducting Qubit Osprey [69] All-to-All $433$ ID, RZ , SX, X Echoed cross resonance gate (ECR) Nov.2022
Eagle [136] $127$ Nov.2021
@@
IonQ Trapped Ions Forte [137] All-to-All $32$ Gpi, GPi2, Virtual Z Fully Entangling MS, Partially Entangling MS 2022
Aria [133] $21$ 2022
Harmony [133] $9$ 2020
Xanadu Photonic Qubits Borealis [134] 3D connectivity $216$ NA NA Jun.2022
X-Series [138] NA NA Mar.2021
Google Superconducting Qubits Sycamore (Weber QC) [63] Square-Lattice Grid $54$ Phased XZ, Virtual Z, Physical Z, Pauli Gates XYZ, PhasedXPowGate Sycamore Gate, Square Root of iSWAP, CZ, FSim Gateset, Wait Gate, Parameterized Gates Oct.2019
Bristlecone [139] $9$-qubit linear array $72$ NA NA Mar.2018
Foxtail [140] NA 2016
Rigetti Superconducting Qubits Aspen-M-3 [141] Octagonal with $3$-fold connectivity ($2$-fold for edges) $80$ RX (angle, qubit) RZ (angle, qubit) CPHASE, MEASURE (qubit, classical reg), CZ Dec.2022
@@
Intel Spin Qubits Intel Quantum Hardware [142] NA $12$ NA NA 2023
Quantinuum (Honeywell) Trapped Ions H2-1 [143] fully connected qubits $32$ Rotation Gates ZZ Aug.2023
@@
Atom Computing Neutral Atoms Phoenix [144] $8$:$1$ connectivity $100$ NA NA Jul.2021
QuEra Neutral Atoms Aquila [145] NA $256$ Rabi Oscillations Time-dependent protocols Dynamic Decoupling Protocols Adiabatic State Preparation and Rydberg Blockade Rabi Frequency Enhancement Levin-Pichler gate analogues Non-Equlibrium dynamics of $2$ qubits Jul.2021
D-Wave Systems Superconducting Qubits Advantage [135] Degree-$15$ Lattice ($15\times 15\times 12$) $5,000$+ superconducting quantum annealing superconducting quantum annealing Sep.2022
2000Q [146] Degree-$6$ Lattice ($8\times 8\times 8$) $2,000$+ Quantum Annealing Quantum Annealing Mar.2017
VI-B1 DiVincenzo’s Criteria
Introduced by theoretical physicist David P. DiVincenzo in the year 2000, DiVincenzo’s criteria [147] defines a set of conditions necessary for constructing a quantum computer. The community widely accepts it, and, in the current state of the field, it represents a useful tool for designing qubits. The following list contains the fundamental characteristics for designing new qubit modalities in suitable experimental setup.
@@ VI-B1 DiVincenzo’s Criteria
VI-B2 Superconducting Qubits
Superconducting circuits are macroscopic in size but detain generic quantum properties such as entanglement, quantized energy levels, and superposition of states, all of which are more commonly associated with atoms. Their quantum state is manipulated using electromagnetic pulses to control the electric charge, the magnetic flux, or the phase difference across a Josephson junction [148] (a device with nonlinear inductance and no energy dissipation). As such, superconducting qubits represent the primitive building blocks of quantum computers [149]. There exist multiple types of superconducting qubits, such as flux qubits, phase qubits, charge qubits, and Transmon qubits. The most popular type is the Transmon qubit [150] given its design capability to limit noise effects, supporting error mitigation and correction.
VI-B3 Semiconducting Quantum Dots & Spin Qubits
Semiconductor charge and spin qubits based on gate-controlled semiconductor quantum dots, shallow dopants (i.e., traces of an impurity element introduced into a chemical material to alter its original electrical or optical properties), and color centers in wide-bandgap materials. Semiconductor Quantum Dots (QDs) [151] are nanoscale material clusters composed of 102102102102105105105105 atoms. The size of the QDs is orders of magnitude larger than a typical atomic radius, yet small enough to provide quantum confinement of electrons and holes in all three spatial dimensions. Depending on their configuration, they can be controlled by both magnetic and electric fields. Therefore, they can be dephased by electric and magnetic field noise, with different types of spin qubits having various control mechanisms and noise susceptibilities.
@@ VI-B3 Semiconducting Quantum Dots & Spin Qubits
VI-B4 Trapped Ions
As the name suggests, the qubits are ions trapped by electric fields and manipulated with lasers [54]. Trapped ions have relatively long coherence times, which implies that the qubits are long-lived. Moreover, they can easily interact with their neighbors.
VI-B5 Neutral Atoms
In the neutral atom quantum computing model, light beams manipulate arrays of single neutral atoms to encode and measure quantum states. In these quantum processors, a qubit is defined by one of two electronic states of an atom, and these single neutral atoms are arranged in configurable arrays. Neutral atom hyperfine qubits provide inherent scalability owing to their identical characteristics, long coherence times, and the ability to be trapped in dense, multidimensional arrays [152]. Combined with the strong entangling interactions provided by Rydberg states [153], all the necessary characteristics for quantum computation are met.
VI-B6 Photonic Atoms
One way to get scalable structures is to use photons (i.e., light particles) to represent qubits [154]. It consists of a ring that is used for photon storage, along with the scattering unit. Unlike other physical systems, photonics allows us access to an infinite number of states.
VI-B7 Transition Metal Dichalcogenides
Among the possibilities in the solid state, a defect in Transition Metal Dichalcogenides (TMDs), known as the nitrogen-vacancy (NV-1) center, stands out for its robustness. Its quantum state may be initialized, manipulated, and measured with high fidelity at room temperature. These controls can be used in conjunction with electronic structure theory to smartly sort through candidate defect systems for probable qubit representation [155]. However, this technology has not really been popular in the industry until now.
VI-B8 Topological Nanowire Qubits (Majorana Qubits)
Majorana particles are a unique class of topological nanowire particles that are their own antiparticles. The quasi-particles emerge from the topological nature of systems and are bound at zero energy. They are predicted to obey non-abelian braiding statistics such that they can “remember” their histories. A pair of Majorana fermions represents a single qubit, and its realization could lead the way to fault-tolerant quantum computing [156]. Due to their stability at close to normal temperatures, Majorana Qubits work fine and are the technology chosen by Microsoft [157].
TABLE IV: Quantum simulators with key features.

### v3

@@ 5.5 Quantum Datasets Collection
Table 2 shows the collection of quantum datasets with key features. The following list provides a brief description of each dataset.
@@
  • The QDataSet (Perrier et al., 2022), released in 2022, is composed of $52$ datasets based on simulations of one-qubit and two-qubit systems evolving in the presence or absence of noise. It provides a large-scale set of datasets for QML practitioners to train, benchmark, and develop classical and quantum algorithms for common tasks in quantum sciences, such as quantum control, quantum tomography, and noise spectroscopy. It has been generated using customized code running on base-level Python packages to facilitate interoperability and portability across common QML platforms. Each dataset consists of $10,000$ samples and includes a range of information (stored in list, matrix, or tensor format) regarding quantum systems and their evolution, such as quantum state vectors, drift and control Hamiltonians, unitaries, Pauli measurement distributions, time series data, pulse sequence data for square and Gaussian pulses, noise and distortion data. The total compressed size of the QDataSet (compressed with Pickle and zip) is around $14$ TB, while its uncompressed size is around $100$ TB.
@@
  • The QM7, QM8, and QM9 datasets are subsets of the GDB-13 (Blum & Reymond, 2009) and GDB-17 (Ruddigkeit et al., 2012) databases, which contain billions of organic molecules. The QM7 (Rupp et al., 2012) is composed of $7,165$ molecules of up to $23$ atoms (including $7$ heavy atoms such as C, N, O, and S), represented using the Coulomb matrix. An extension of the QM7 dataset for multitask learning, called QM7B (Montavon et al., 2013), has $13$ additional properties (e.g., polarizability, HOMO and LUMO eigenvalues, excitation energies) and contains a total of $7,211$ molecules. The QMB8 (Ramakrishnan et al., 2015) has a training set of $10,000$ molecules with up to $8$ heavy atoms. The QM9 (Ramakrishnan et al., 2014) contains $133,885$ molecules with up to $9$ heavy atoms. These databases serve as benchmarks for hybrid QML models and systematic identification of structure-property relationships.
@@
  • The PennyLane Datasets (Bergholm et al., 2018) are composed of the Quantum Chemistry (QChem) Datasets for common molecular systems and the Quantum Many-Body Physics (QMBP) Datasets. The QChem datasets contain the electronic structure data for $42$ different geometries of molecules such as linear hydrogen chains, metallic and non-metallic hydrides, and charged species. For each geometry of the molecule, it is possible to extract molecular data (i.e., information regarding the molecule), Hamiltonian data (i.e., Hamiltonian for the molecular system), tapering data (i.e., features based on Z2 symmetries of the molecular Hamiltonian for performing tapering), tapered observables data, and variational data. The QMBP datasets contain spin models displaying quantum correlations. For each spin system, datasets are generated for $1$-D lattices (linear chain) and $2$-D lattices (rectangular grid). Each dataset contains values for $100$ different configurations of tunable parameters such as the external magnetic field and coupling constants.
@@
  • The TensorFlow Quantum Data builds upon classical datasets such as MNIST (LeCun et al., 1998) and Fashion-MNIST (Xiao et al., 2017). It implements the method proposed by Farhi & Neven (2018) to use binary encoding for representing each pixel with a qubit, with the state depending on the pixel’s value. According to the experimental results of Huang et al. (2021), this method achieves quantum speedup in the fault-tolerant regime compared to the classical system.
@@
  • The Mendley Datasets for Quantum Circuits Mapping (Acampora et al., 2021) helps to solve the circuit mapping problem as a classification task. Each dataset contains features related to the calibration data of the physical device and others related to the generated quantum circuit for specific IBM quantum machines, such as IBMQ Santiago, IBMQ Athens, and IBMQ 16 Melbourne.
@@
  • The QFlow Lite dataset (Zwolak et al., 2018) contains a Python-based software suite to train neural networks to recognize the state of a device and differentiate between states in experimental data. It consists of $1,001$ idealized simulated measurements with gate configurations sampling (stored as $100\times 100$ pixel maps) over different realizations of the same type of device. The QFlow Lite represents a reference dataset for researchers to use in their experiments for developing QML approaches and concepts. The expanded dataset, denominated QFlow 2.0, consists of $1,599$ idealized simulated measurements stored as $250\times 250$ pixel maps. In addition, the QFlow 2.0 dataset includes two sets of noisy simulated measurements, one with a noise level varied around $1.5$ times the optimized noise level, and the other with a noise level ranging from 00 to $7$ times the optimized noise level.
@@
  • The NTangled Datasets (Schatzki et al., 2021) consist of trained weights for three hardware-efficient Ansatzes), strongly entangling, and convolutional, varying the number of qubits, depths, and types of multipartite entanglement. The states are generated by sending product state inputs into PQCs. The NTangled Datasets are used to benchmark QML models for supervised learning classification tasks.
6 Tools and Technologies
@@
6.1 Quantum Computing Models
@@
The first concept of a quantum computer was introduced in 1982 by Feynmann (1982). Since then, a significant amount of research efforts have been conducted that resulted in the development of real quantum computers at our disposal for experiments (IBM, b). Due to their different use cases and architectures, quantum computers can be classified into quantum gate models (or quantum circuits) and quantum annealers.
6.1.1 Quantum Gate Model
@@
The gate model involves operations in quantum computation using matrix operators, which are applied to the input qubits. These operators process the information and generate the final output distributions. They are designed with unitary constraints to qualify as valid quantum operators, making them reversible. The reversibility property is necessary for introducing some form of determinism in the quantum computation. A single operator is referred to as a gate in the quantum circuit and can be applied independently or grouped together with other similar quantum operators to perform computation on qubits. A gate can be combined with other gates in varying orientations and series such that they solve certain problems. The collection of gates arranged in configurations respective to the problem being solved along with qubits and measurements is referred to as a circuit (see Section 2.4). Complex problems in the quantum domain can be solved by configuring gates in meaningful ways. The computation done with this quantum model is referred to as the gate model. Since it is easily interpretative and intuitive, it is nowadays the most used quantum computation model.
6.1.2 Quantum Annealer
@@
Quantum Annealing (QA) is an adiabatic quantum computation method that finds optimal solutions to problems. It uses quantum tunneling, entanglement, and superposition to find the solution in terms of the system’s energy. It usually corresponds to finding the global energy minimum in the quantum energy state plane. Instead of using temperature to explore the problem space, QA uses the laws of quantum mechanics to measure the energy state. A quantum annealer is relatively more noise-tolerant compared to the gate model. Quantum annealers shine when they are used to search for the best from many possible combinations. Such challenges are called combinatorial optimization problems (see Section 4.3.4).
6.2 Quantum Hardware
@@
Single qubits can be implemented using different materials and principles. As shown in Figure 32, companies do not use the same technology for implementing their quantum computers. For example, IBM Quantum uses niobium and aluminum on a silicon base to achieve superconductivity under near absolute zero temperature (IBM, 2022). IonQ uses trapped ion quantum made of Ytterbium, a rare metal (IonQ, ). Another example is Xanadu, which applies photonics (light particles) in the quantum processing unit (Madsen et al., 2022). Using ion and light has advantages over the superconductivity materials as it works under room temperature. The superconducting qubits are more scalable and are implemented by other companies like Google (Frank Arute, 2019) and D-Wave (McGeoch & Farre, 2020). It is exciting to see future innovations in this area. In this section, after discussing the set of conditions necessary for designing qubits (denoted as DiVincenzo’s criteria), we review the qubit technologies to understand their advantages and disadvantages and the reasons for their adoption in the industry. A detailed view of the quantum hardware technologies is also illustrated in Table 3.
@@
Table 3: Quantum hardware technologies with key features.
@@
IBM Superconducting Qubit Osprey (IBM, 2022) All-to-All $433$ ID, RZ , SX, X Echoed cross resonance gate (ECR) Nov.2022
Eagle (IBM, a) $127$ Nov.2021
@@
IonQ Trapped Ions Forte (Chen et al., 2023) All-to-All $32$ Gpi, GPi2, Virtual Z Fully Entangling MS, Partially Entangling MS 2022
Aria (IonQ, ) $21$ 2022
Harmony (IonQ, ) $9$ 2020
Xanadu Photonic Qubits Borealis (Madsen et al., 2022) 3D connectivity $216$ NA NA Jun.2022
X-Series (Arrazola et al., 2021) NA NA Mar.2021
Google Superconducting Qubits Sycamore (Weber QC) (Frank Arute, 2019) Square-Lattice Grid $54$ Phased XZ, Virtual Z, Physical Z, Pauli Gates XYZ, PhasedXPowGate Sycamore Gate, Square Root of iSWAP, CZ, FSim Gateset, Wait Gate, Parameterized Gates Oct.2019
Bristlecone (Google, 2018) $9$-qubit linear array $72$ NA NA Mar.2018
Foxtail (Goo, 2016) NA 2016
Rigetti Superconducting Qubits Aspen-M-3 (Computing, ) Octagonal with $3$-fold connectivity ($2$-fold for edges) $80$ RX (angle, qubit) RZ (angle, qubit) CPHASE, MEASURE (qubit, classical reg), CZ Dec.2022
@@
Intel Spin Qubits Intel Quantum Hardware (Intel, ) NA $12$ NA NA 2023
Quantinuum (Honeywell) Trapped Ions H2-1 (Quantinuum, ) fully connected qubits $32$ Rotation Gates ZZ Aug.2023
@@
Atom Computing Neutral Atoms Phoenix (Barnes et al., 2022) $8$:$1$ connectivity $100$ NA NA Jul.2021
QuEra Neutral Atoms Aquila (Wurtz et al., 2023) NA $256$ Rabi Oscillations Time-dependent protocols Dynamic Decoupling Protocols Adiabatic State Preparation and Rydberg Blockade Rabi Frequency Enhancement Levin-Pichler gate analogues Non-Equlibrium dynamics of $2$ qubits Jul.2021
D-Wave Systems Superconducting Qubits Advantage (McGeoch & Farre, 2020) Degree-$15$ Lattice ($15\times 15\times 12$) $5,000$+ superconducting quantum annealing superconducting quantum annealing Sep.2022
2000Q (King et al., 2017) Degree-$6$ Lattice ($8\times 8\times 8$) $2,000$+ Quantum Annealing Quantum Annealing Mar.2017
6.2.1 DiVincenzo’s Criteria
Introduced by theoretical physicist David P. DiVincenzo in the year 2000, DiVincenzo’s criteria (DiVincenzo, 2000) defines a set of conditions necessary for constructing a quantum computer. The community widely accepts it, and, in the current state of the field, it represents a useful tool for designing qubits. The following list contains the fundamental characteristics for designing new qubit modalities in suitable experimental setup.
@@ 6.2.1 DiVincenzo’s Criteria
6.2.2 Superconducting Qubits
Superconducting circuits are macroscopic in size but detain generic quantum properties such as entanglement, quantized energy levels, and superposition of states, all of which are more commonly associated with atoms. Their quantum state is manipulated using electromagnetic pulses to control the electric charge, the magnetic flux, or the phase difference across a Josephson junction (Kockum & Nori, 2019) (a device with nonlinear inductance and no energy dissipation). As such, superconducting qubits represent the primitive building blocks of quantum computers (Bravyi et al., 2022). There exist multiple types of superconducting qubits, such as flux qubits, phase qubits, charge qubits, and Transmon qubits. The most popular type is the Transmon qubit (Salmanogli, 2021) given its design capability to limit noise effects, supporting error mitigation and correction.
6.2.3 Semiconducting Quantum Dots & Spin Qubits
Semiconductor charge and spin qubits based on gate-controlled semiconductor quantum dots, shallow dopants (i.e., traces of an impurity element introduced into a chemical material to alter its original electrical or optical properties), and color centers in wide-bandgap materials. Semiconductor Quantum Dots (QDs) (Harvey, 2022) are nanoscale material clusters composed of 102102102102105105105105 atoms. The size of the QDs is orders of magnitude larger than a typical atomic radius, yet small enough to provide quantum confinement of electrons and holes in all three spatial dimensions. Depending on their configuration, they can be controlled by both magnetic and electric fields. Therefore, they can be dephased by electric and magnetic field noise, with different types of spin qubits having various control mechanisms and noise susceptibilities.
@@ 6.2.3 Semiconducting Quantum Dots & Spin Qubits
6.2.4 Trapped Ions
As the name suggests, the qubits are ions trapped by electric fields and manipulated with lasers (Nielsen & Chuang, 2010a). Trapped ions have relatively long coherence times, which implies that the qubits are long-lived. Moreover, they can easily interact with their neighbors.
6.2.5 Neutral Atoms
In the neutral atom quantum computing model, light beams manipulate arrays of single neutral atoms to encode and measure quantum states. In these quantum processors, a qubit is defined by one of two electronic states of an atom, and these single neutral atoms are arranged in configurable arrays. Neutral atom hyperfine qubits provide inherent scalability owing to their identical characteristics, long coherence times, and the ability to be trapped in dense, multidimensional arrays (Henriet et al., 2020). Combined with the strong entangling interactions provided by Rydberg states (Shi, 2022), all the necessary characteristics for quantum computation are met.
6.2.6 Photonic Atoms
One way to get scalable structures is to use photons (i.e., light particles) to represent qubits (Slussarenko & Pryde, 2019). It consists of a ring that is used for photon storage, along with the scattering unit. Unlike other physical systems, photonics allows us access to an infinite number of states.
6.2.7 Transition Metal Dichalcogenides
Among the possibilities in the solid state, a defect in Transition Metal Dichalcogenides (TMDs), known as the nitrogen-vacancy (NV-1) center, stands out for its robustness. Its quantum state may be initialized, manipulated, and measured with high fidelity at room temperature. These controls can be used in conjunction with electronic structure theory to smartly sort through candidate defect systems for probable qubit representation (Tsai et al., 2022). However, this technology has not really been popular in the industry until now.
6.2.8 Topological Nanowire Qubits (Majorana Qubits)
Majorana particles are a unique class of topological nanowire particles that are their own antiparticles. The quasi-particles emerge from the topological nature of systems and are bound at zero energy. They are predicted to obey non-abelian braiding statistics such that they can “remember” their histories. A pair of Majorana fermions represents a single qubit, and its realization could lead the way to fault-tolerant quantum computing (Tutschku et al., 2020). Due to their stability at close to normal temperatures, Majorana Qubits work fine and are the technology chosen by Microsoft (Aghaee et al., 2023).
Table 4: Quantum simulators with key features.